OpenStudy (anonymous):

The half-life for the radioactive decay of calcium-47 is 4.5days. If a sample has an activity of 3.0Ci after 9.0days, what was the initial activity of the sample?

OpenStudy (xishem):

Well, think about it logically. If it's been decaying for 9.0 days and its half-life is 4.5 days, how many half-lifes has it gone through?

OpenStudy (xishem):

In a more analytic sense... This situation can be modeled by this equation...\[N(t)=N_0\frac{1}{2}^{\frac{t}{t_{1/2}}}\]Where... N(t) is the amount (or activity in this case) at time t as a function of time N_0 is the initial amount (activity) t is the time t_{1/2} is the half-life So...\[3.0Ci=N_0*(\frac{1}{2})^\frac{9.0}{4.5}\]\[3.0Ci=N_0*(1/2)^2 \rightarrow 3.0Ci = N_0*0.25\]\[N_0=12.0Ci \]